OpenStudy (anonymous):

A straight line passes through the points (3, −9) and (−5, 1). Write an equation for this line. a. 8y + 10x = −44 b. −8y + 10x = −42 c. −8y + 10x = −32 d. 8y + 10x = −42 e. 8y + 10x = 58

OpenStudy (anonymous):

You can get the slope first, like we did in your last question. Then, you can do: y - y1 = m(x - x1) where x1 and y1 are any specific point that you know (you know of two points, so just pick one). Then rearrange.

OpenStudy (anonymous):

how do you rearrange?

OpenStudy (anonymous):

Just by putting it in the form that all the choices are in, which is variables on the left and a number on the right.

OpenStudy (anonymous):

So, that first step is getting the slope like in the previous problem. Remember, it's \[(y _{1} - y _{2}) / (x _{1} - x _{2}) = m\] which is the slope.

OpenStudy (anonymous):

For the second step, take the m which you calculated in step 1, and put it in the equation I stated in the first post above. Take a point (either will do) for x1 and y1. Plug those in also. Multiply out. Then put the variables (with their numeric coefficients) over on the left and the leftover number over to the right.

OpenStudy (anonymous):

the slope is -11/8?

OpenStudy (anonymous):

You're slightly off. I think if you check your arithmetic, you'll get it. You're very close.

OpenStudy (anonymous):

[1 - (-9)] / [-5 - 3] for slope.

OpenStudy (anonymous):

10/8

OpenStudy (anonymous):

Not quite. You're having trouble with arithmetic. [1 - (-9)] / [-5 - 3] = 10 / (-8). Go over the subtraction which is done separately in the numerator and the denominator.

OpenStudy (anonymous):

Oh sorry, I missed the negative right there.

OpenStudy (anonymous):

So, slope is -10 / 8. Now, go to my post #4 directions (which are an embellishment of my first post).

OpenStudy (anonymous):

y-(-9)=-10/8(x-3)?

OpenStudy (anonymous):

Yes, if you mean y-(-9)=(-10/8)(x-3)? I added some parentheses.

OpenStudy (anonymous):

Oh yeah, that's what I mean.

OpenStudy (anonymous):

Some simplification steps are y-(-9) = y+9 and it would be good to multiply both sides by 8 to get rid of the 8 in the first factor on the right side.

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

Are you all set now, because I have to leave soon. Any last questions?

OpenStudy (anonymous):

8y + 10x = −42, thanks for helping. sorry for all the hassle.

OpenStudy (anonymous):

np No hassle. You're a good hard worker! And you got it right!